# Paying and Persuading

### v4

@@ 6 Optimal Loyalty Contracts
Return now to the repeated interaction between Uber and a driver. Suppose each period (i.e. a day) Uber pushes $n$ rides to a driver, the quality of which is unknown to the driver but privately observed by Uber. Suppose that Uber always wants the driver to accept the ride ($a_{i}=1$) (though their value from acceptance may vary from ride to ride), so that the payoff from ride $i$ is given by $v_{i}(a_{i},\theta_{i})=c_{i}\mathbf{1}\{a_{i}=1\}$ for some $c_{i}>0$. The driver, however, wants to accept the ride if and only if the ride is good, $\theta_{i}=1$; normalize their payoff so that $u_{i}(a_{i},\theta_{i})=\mathbf{1}\{\theta_{i}=a_{i}\}$. [^18] In each dimension, this describes the judge-prosecutor game a la Kamenica and Gentzkow (2011) where Uber is the prosecutor. An alternative (less natural) interpretation of this setup is one where the prosecutor tries many simultaneous trials in front of a judge in each period, and can also pay the judge to make a decision (though as $k\to\infty$, I recover the case without payments). To incentivize the driver to accept rides, Uber can use two tools: directly increase their payments, or engage in information design.
@@
This formulation encodes three assumptions. First, that payoffs across dimensions are additively separable. Second, that the efficiency of transfers $k$ (implicit in the model setup) is the same across dimensions. Third, that the driver’s prior belief about the quality of the ride is independently drawn across periods according to some prior belief $\mu_{0}^{i}$ in dimension $i$, with $\mu_{0}=\prod_{i}\mu_{0}^{i}$. Suppose moreover that $\theta_{i}\sim\mu_{0}^{i}$ across all time periods, so that ride $i$’s valuation is drawn i.i.d. across time, and that $\mu_{0}^{i}\in(0,\frac{1}{2})$ for each $i$, so that the problem is nontrivial in each dimension.
@@
A tiered loyalty contract is a strategy $\sigma$ and a sequence of increasing stopping times $\{T_{i}^{*}\}_{i=1}^{n}$, $T_{i}^{*}\leq T_{i+1}^{*}$ such that for $\mathbb{P}^{\sigma}$-almost all histories,
$\rho_{i}(h^{\xi};\sigma)=\begin{cases}\rho_{i}^{FI}\text{ if }t>T_{i}^{*}(h)\\ \rho_{i}^{BP}\text{ if }t<T_{i}^{*}(h)\end{cases}.$
Tiered loyalty contracts allow for one period of indeterminacy ($T_{i}^{*}(h)=T$), which arises due to the integer programming problem when providing incentives. In general, though tiered loyalty programs are stark: they provide myopically optimal information up until a certain point at which the relationship matures and Sender’s “stock” of future informational incentives runs out. After this, Sender “promotes” the driver, gives them full information, and then pays them to take valuable rides. This transition happens “dimension-by-dimension:” Sender slowly unwinds their advantage by transitioning from the static optima directly to full information in each dimension separately; consequently, each transition $T_{i}^{*}$ can be thought of as a distinct “tier” in the relationship.
@@
Proposition 10.
Suppose $c_{i}>k$ for some $i$. There is an optimal strategy which features transfers as a last resort and is a tiered loyalty contract.
The proof can be found in PROOF OF PROPOSITION 10 The approximate intuition is as follows. First, the Pareto frontier is governed by the summary statistic $\mathbb{P}(a_{i}=1|\theta_{i}=0)$ at each time in each dimension, the rate at which persuasion is “successful.” For Sender to provide incentives via information, it must be that they commit to an information structure which decreases this probability, which gives Receiver a marginal unit of 1-1 while costing Sender $c_{i}$. So long as this probability is interior (between the static optimum and full information), the Pareto frontier is linear, and hence dynamic incentives can be “backloaded” using a similar squeezing procedure as Lemma 5 to provide information in dimension $i$ tomorrow instead of today. Hence it is possible to provide incentives by promising full information in some dimension at some point in the future, until eventually the informational advantage is completely unwoven; at this point, Sender unwinds their entire informational advantage and transfers come in as a last resort to motivate the agent. The assumption there is a dimension where $c_{i}>k$ finally guarantees that in each period, Sender wishes to provide intertemporal incentives to motivate Receiver to take action $a_{i}=1$ even when the state is bad.
Proposition 10 mirrors the loyalty contracts employed by Uber, who run a driver loyalty program with four tiers (“Blue,” “Gold,” “Platinum,” and “Diamond”) each of which have different benefits and differential access to information. I thus see my model as providing a potential explanation for Uber’s driver loyalty program. Alternatively, with two projects (like the project selection example succeeding Proposition 8 without a safe outside option), the time $\xi_{i}^{*}$ can be thought of as promotion: once the worker has chosen the “bad-fit” project enough times, the manager gives them information (and thus discretion) to choose projects which the employee would prefer, even if doing so runs against the manager’s own interests.
@@ 7 Discussion
This paper studied a model of dynamic contracting where Sender can both pay and persuade as ways to motivate Receiver. Theorem 1 characterizes the time structure of optimal transfers—payments occur on the path of play only if Sender has committed already to reveal the state forever afterwards. One interpretation of this result is that dynamic information endows the Sender with a “stock” of informational incentives, which they can always draw down upon before turning to transfers. In simple settings—such as the Uber example—this observation implied to the optimality of tiered loyalty programs, where static persuasion remains optimal until the entire informational stock is exhausted, after which Sender fully reveals the state and simply pays Receiver.
There are several natural extensions under which I conjecture the result continues to hold. First, a modified Theorem 1 extends to any restricted set of experiments where there is a unique Blackwell maximal element. My backloading result may also extend to the case where transfers are nonlinear in cost but differentiable. Second, the result extends to cases where payoffs evolve in every period, and the state is noisily observed, as discussed after Theorem 1.
Several questions about the model remain open and are fruitful areas for future research. First, information was free to generate in my model; I conjecture for suitably defined costs of information (such that full information is not infinitely costly), an adapted version of the result holds. Second, understanding the model with moral hazard seems like a natural enrichment of the model. Importantly, Lemma 4 may no longer apply, as the action is not perfectly monitored and thus the correct “pullback utility” at the exact optimum is no longer clear. Despite this, tools from the quota mechanisms literature could prove effective in characterizing approximately optimal contracts where payments are a last resort. Third, it is natural to consider the case with multiple Receivers, who are playing some type of coordination game against Sender (i.e. a dynamic version of Mathevet et al. (2020) and Smolin and Yamashita (2025) with Winter (2004) and Halac et al. (2021). Finally, characterizing explicitly dynamic contracts outside of the simple cases studied in the previous section remains an open question. Indeed, Lehrer and Shaiderman (2025) call the simpler version of this question, without transfers, the “grand question” of dynamic information design.
@@
Fix any obedient strategy. At each on-path history, replace the signal $s$ with a distribution of random signals $\{(\theta|s,a)\}_{a\in\text{supp}(\alpha)}$ where the probability of each signal being sent is exactly $\alpha(a)$. Clearly this does not affect Receiver’s information or their continuation incentives and hence the realized belief is $\theta|s$ and the recommended action is obeyed. Second, note decreasing transfers as much as possible at actions not induced on path can only weaken incentive compatibility constraints and help support a conjectured equilibrium profile. ∎
@@
The first part is standard, noting that the maximizing set of the static problem is nonempty and that payoffs satisfy a transversality condition (by (BD)). The second follows if $u^{\prime}(\mu,a)=\mathcal{U}(\{h^{\xi},\mu,a\})$, then the choice of future utility promise and payoffs today must be part of an optimum to the sequential formulation, as otherwise a profitable one-shot deviation exists in the sequential problem. ∎
@@
The first statement. For any $(u,\mu_{0})$ and $\{\rho,t,u^{\prime},\alpha\}\in\mathcal{F}(u,\mu_{0})$, and $\varepsilon>0$, let $\tilde{V}(u+\varepsilon,\mu)$ be the value for Sender by choosing at inputs $(u+\varepsilon,\mu_{0})$ the (feasible) tuple $\{\rho,t,u^{\prime}+\frac{\varepsilon}{\delta},\alpha\}$. This tuple strictly increases Receiver utility to meet the higher utility promise without otherwise affecting any IC constraints, so in particular it must do weakly worse than the optimum: $V(u+\varepsilon,\mu_{0})\geq\tilde{V}(u+\varepsilon,\mu_{0})$.
Suppose now the lemma is false, so $V_{+}^{\prime}(u^{\prime}(\mu,a),\mu)>-k$ for some $u^{\prime}(\mu,a)$ where $\mu\in\text{supp}(\rho)$, $a\in\text{supp}(\alpha)$. By construction, we have
@@
$\tilde{V}(u+\varepsilon,\mu_{0})-V(u,\mu_{0})=\delta\mathbb{E}_{\rho,\alpha}\left[V\left(u^{\prime}(\mu,a)+\frac{\varepsilon}{\delta},\mu\right)-V(u^{\prime}(\mu,a),\mu)\right].$
@@
$=\lim\limits_{\varepsilon\to 0_{+}}\frac{\delta}{\varepsilon}\mathbb{E}_{\rho,\alpha}\left[V\left(u^{\prime}(\mu,a)+\frac{\varepsilon}{\delta},\mu\right)-V(u^{\prime}(\mu,a),\mu)\right]$
$=\mathbb{E}_{\rho,\alpha}\left[V_{+}^{\prime}(u^{\prime}(\mu,a),\mu)\right]>-k$
where the second equality interchanges the (finite) expectation and evaluates the right derivative explicitly. The final inequality uses the fact that $V_{+}^{\prime}(\cdot,\mu)\geq-k$ everywhere with strict inequality (by assumption) for at least one $(\mu,a)\in\text{supp}(\rho,\alpha)$. This contradicts our assumption about $V_{+}^{\prime}(u,\mu_{0})$.
The second statement. Suppose not. Then we can find tuple $\{\rho,t,u^{\prime},\alpha\}\in\mathcal{F}(u,\mu_{0})$ at some $(u,\mu_{0})$ where $t(\mu,a)>0$, $\mu\in\text{supp}(\rho)$ $a\in\text{supp}(\alpha)$, but $V_{+}^{\prime}(u^{\prime}(\mu,a),\mu)>-k$. Recall Receiver’s incentive compatibility constraint when recommended $a$ at $\mu$ takes the form
@@
Consider now the alternative compensation scheme $\{\rho,\tilde{t},\tilde{u},a\}$ which leaves the information and action recommendations unchanged by changing the compensation scheme only at $\mu$: for some $\varepsilon>0$ (restricted to be small enough that $\tilde{t}(\mu)$ remains positive), set $\tilde{t}(\mu,a)=t(\mu,a)-\varepsilon$, and set $\tilde{u}(\mu,a)=u(\mu,a)+\frac{1-\delta}{\delta}\varepsilon$ at all other beliefs $\tilde{t}=t$ and $\tilde{u}=u$. Then by construction
$(1-\delta)\tilde{t}(\mu)+\delta\tilde{u}=(1-\delta)t(\mu)+\delta u$
which implies Receiver’s incentive compatibility constraints are unperturbed under $\{\rho,\tilde{t},\tilde{u},a\}$ (as is their promise-keeping constraint). Hence, $\{\rho,\tilde{t},\tilde{u},a\}$ is feasible at $(u,\mu_{0})$. Whenever belief $\mu$ is realized and $a$ is recommended, however, Sender’s payoff differs by a value of
$(1-\delta)(\tilde{t}(\mu,a)-t(\mu,a))+\delta(V(\tilde{u}(\mu,a),\mu)-V(u(\mu,a),\mu))$
$=(1-\delta)k\varepsilon+\left[\delta V\left(u(\mu,a)+\frac{1-\delta}{\delta}\varepsilon,\mu\right)-\delta V(u(\mu,a),\mu)\right].$
@@
$\lim\limits_{\varepsilon\to 0_{+}}\left((1-\delta)k+\delta\frac{V\left(u(\mu,a)+\frac{1-\delta}{\delta}\varepsilon,\mu\right)-V(u(\mu,a),\mu)}{\varepsilon}\right)=(1-\delta)k+(1-\delta)V_{+}^{\prime}(u(\mu,a)).$
@@
Start from $u$-constrained optimal $\sigma$ and define $\bar{\sigma}$ as follows. Let $\rho^{FI}$ be the experiment that reveals the state, and suppose $\bar{\rho}(h^{\xi};\bar{\sigma})=\rho^{FI}$ always. Suppose Sender recommends some action for sure $\bar{\alpha}(\delta_{\theta};\bar{\sigma})=\delta_{a(\theta)}$, where $a(\theta)\in\operatorname*{\arg\!\max}_{a\in A}u(a,\theta)$ is some action that maximizes Receiver’s utility at that state, and which promises Receiver $u(\delta_{\theta},a)=u^{RFI}(\delta_{\theta})$ in the future at all recommended actions (and $\underline{U}(\delta_{\theta})$ otherwise). Suppose moreover that payments never occur on the path of play, $\bar{t}\equiv 0$. Clearly $\{\bar{\rho},\bar{t},\bar{\alpha}\}$ is a feasible strategy at any $u\leq u^{RFI}(\mu_{0})$ and gives Receiver a payoff of exactly $u^{RFI}(\mu_{0})$.
Now define the following strategy $\sigma^{\prime}$. Still always reveal the state at each history, and let $\alpha^{\prime}$ be defined as follows: for each state $\theta$ and on-path $h^{\xi}$, set $\alpha^{\prime}(h^{\xi};\sigma^{\prime})(\delta_{\theta})=\mathbb{Q}_{\xi}^{\sigma}(\theta)$ and let transfers $t^{\prime}$ be such that for each $a^{\prime}\in\text{supp}(\alpha^{\prime}(h^{\xi};\sigma^{\prime})(\delta_{\theta}))$,
@@
$(1-\delta)\max_{a\in A}u(a,\theta)+\delta u^{RFI}(M\delta_{\theta})\geq(1-\delta)u(\tilde{a},\theta)+\underline{U}(M\delta_{\theta})$
@@
We thus now have a profile $\{\bar{\rho},t^{\prime},\alpha^{\prime}\}$ that, by construction, satisfies $\mathbb{Q}_{\xi}^{\sigma}=\mathbb{Q}_{\xi}^{\sigma^{\prime}}$ at every possible history and where the state is fully revealed, i.e. properties (1) and (2). Moreover, it gives (by construction) Sender the same expected payoff absent transfers. How much must Sender pay Receiver in expectation under $\sigma^{\prime}$? Note by construction payments are exactly equal to
$(1-\delta)\sum_{\xi=0}^{\infty}\delta^{\xi}\mathbb{E}_{Q_{\xi}^{\sigma}}\left[\max_{a\in A}u(a,\theta)-u(a^{\prime},\theta)\right]=u^{RFI}(\mu_{0})-\mathcal{U}(\sigma),$
where we use the fact
@@
$\mathcal{V}(\sigma^{\prime})=\mathcal{V}(\sigma)-k\left(u^{RFI}(\mu_{0})-\mathcal{U}(\sigma)\right)$

### v5

@@ 6 Optimal Loyalty Contracts
Consider the following stylized environment, inspired by the problem Uber must face when designing a dynamic contract. Suppose each period (i.e. a day) Uber (Sender) pushes $n$ rides to a driver (Receiver), the quality of which is unknown to the driver but privately observed by Uber. Suppose that Uber always wants the driver to accept the ride ($a_{i}=1$) (though their value from acceptance may vary from ride to ride), so that the payoff from ride $i$ is given by $v_{i}(a_{i},\theta_{i})=c_{i}\mathbf{1}\{a_{i}=1\}$ for some $c_{i}>0$. The driver, however, wants to accept the ride if and only if the ride is good, $\theta_{i}=1$; normalize their payoff so that $u_{i}(a_{i},\theta_{i})=\mathbf{1}\{\theta_{i}=a_{i}\}$. [^21] In each dimension, this describes the judge-prosecutor game a la Kamenica and Gentzkow (2011) where Uber is the prosecutor. An alternative (less natural) interpretation of this setup is one where the prosecutor tries many simultaneous trials in front of a judge in each period, and can also pay the judge to make a decision (though as $k\to\infty$, I recover the case without payments). To incentivize the driver to accept rides, Uber can use two tools: directly increase their payments, or engage in information design.
@@
This formulation encodes three assumptions. First, that payoffs across dimensions are additively separable. Second, that the efficiency of transfers $k$ (implicit in the model setup) is the same across dimensions. Third, that the driver’s prior belief about the quality of the ride is independently drawn across periods according to some prior belief $\mu_{0}^{i}$ in dimension $i$, with $\mu_{0}=\prod_{i}\mu_{0}^{i}$. Suppose moreover that $\theta_{i}\sim\mu_{0}^{i}$ across all time periods, so that ride $i$’s valuation is drawn i.i.d. across time, and that $\mu_{0}^{i}\in(0,\frac{1}{2})$ for each $i$, so that there is room for incentive provision in each dimension.
@@
A tiered loyalty contract is a contract $\sigma$ and a sequence of increasing stopping times $\{T_{i}^{*}\}_{i=1}^{n}$, $T_{i}^{*}\leq T_{i+1}^{*}$ such that for $\mathbb{P}^{\sigma}$-almost all histories,
$\rho_{i}(h^{\xi};\sigma)=\begin{cases}\rho_{i}^{FI}\text{ if }t>T_{i}^{*}(h^{\xi})\\ \rho_{i}^{BP}\text{ if }t<T_{i}^{*}(h^{\xi})\end{cases}.$
Tiered loyalty contracts allow for one period of indeterminacy ($T_{i}^{*}(h^{\xi})=T$), which arises due to the integer programming problem when providing incentives. In general, though tiered loyalty programs are stark: they provide myopically optimal information up until a certain point at which the relationship matures and Sender’s “stock” of future informational incentives runs out. After this, Sender “promotes” the driver, gives them full information, and then pays them to take valuable rides. This transition happens “dimension-by-dimension:” Sender slowly unwinds their advantage by transitioning from the static optima directly to full information in each dimension separately; consequently, each transition $T_{i}^{*}$ can be thought of as a distinct “tier” in the relationship.
@@
Theorem 2.
There is an optimal contract which features transfers as a last resort and is a tiered loyalty contract.
Theorem 2 is a deuteragonist of this paper. In particular, it explicitly solves for the dynamic optimal contract in a class of payoffs that generalize the leading example in the information design literature. Moreover, thus a contract is natural and has a “bang-bang” structure: myopically optimal information is provided until some random promotion time, after which the state in each dimension is fully revealed. Importantly, even though the problem (and incentives) are separable, information is leveraged across dimensions—future information revelation in one dimension is used as an incentive tool to encourage the agent to take the high action in a different dimension.
Theorem 2 mirrors the loyalty contracts employed by Uber, who run a driver loyalty program with four tiers (“Blue,” “Gold,” “Platinum,” and “Diamond”) each of which have different benefits and differential access to information. I thus see my model as providing a potential explanation for Uber’s driver loyalty program. Alternatively, with two projects (like the project selection example succeeding Proposition 9 without a safe outside option), the time $\xi_{i}^{*}$ can be thought of as promotion: once the worker has chosen the “bad-fit” project enough times, the manager gives them information (and thus discretion) to choose projects which the employee would prefer, even if doing so runs against the manager’s own interests.
The proof can be found in PROOF OF THEOREM 2 The approximate intuition is as follows. First, the Pareto frontier is governed by the summary statistic $\mathbb{P}(a_{i}=1|\theta_{i}=0)$ at each time in each dimension, the rate at which persuasion is “successful.” For Sender to provide incentives via information, it must be that they commit to an information structure which decreases this probability, which gives Receiver a marginal unit of $1$ while costing Sender $c_{i}$. So long as this probability is interior (between the static optimum and full information), the Pareto frontier is linear, and hence dynamic incentives can be “backloaded” using a similar squeezing procedure as Lemma 5 to provide information in dimension $i$ tomorrow instead of today. Hence it is possible to provide incentives by promising full information in some dimension at some point in the future, until eventually the informational advantage is completely unwoven; at this point, Sender unwinds their entire informational advantage and transfers come in as a last resort to motivate the agent.
@@ 7 Discussion
This paper studied a model of dynamic contracting where Sender can both pay and persuade as ways to motivate Receiver. Theorem 1 characterizes the time structure of optimal transfers—there exists an optimal contract where payments occur on the path of play only if Sender has committed already to reveal the state forever afterwards. One interpretation of this result is that dynamic information endows Sender with a “stock” of informational incentives, which they can always draw down upon before turning to transfers. In simple settings—such as the Uber example—this observation implied to the optimality of tiered loyalty programs, where static persuasion remains optimal until the entire informational stock is exhausted, after which Sender fully reveals the state and simply pays Receiver.
There are several natural extensions under which I conjecture the result continues to hold. A modified Theorem 1 extends to any restricted set of experiments where there is a unique Blackwell maximal element. My backloading result may also extend to the case where transfers are nonlinear in cost but differentiable. The result should extend to cases where payoffs evolve over time and the state is noisily observed, as discussed after Theorem 1.
Several questions about the model remain open and are fruitful areas for future research. First, information was free to generate in my model; I conjecture for suitably defined costs of information (such that full information is not infinitely costly), an adapted version of the result holds. Second, understanding the model with moral hazard seems like a natural enrichment of the model. Importantly, Lemma 4 may no longer apply, as the action is not perfectly monitored and thus the correct “pullback utility” at the exact optimum is no longer clear. Despite this, tools from the quota mechanisms literature could prove effective in characterizing approximately optimal contracts where payments are a last resort. Third, it is natural to consider the case with multiple Receivers, who are playing some type of coordination game against Sender (i.e. a dynamic version of Mathevet et al. (2020) and Smolin and Yamashita (2026) with Winter (2004) and Halac et al. (2021). Fourth, comparative statics are an open question. It remains to be seen how the stopping time $T_{\sigma}$ which determines when transfers are used in contracts where transfers are a last resort changes with model primitives, in particular the discount rate. As $\delta$ increases, intertemporal incentives become more attractive to the principal, though the value of future information also increases, so it is not obvious whether transfers get used sooner or later. Finally, characterizing explicitly dynamic contracts outside of the simple cases studied in the previous section remains an open question. Indeed, Lehrer and Shaiderman (2025) call the simpler version of this question, without transfers, the “grand question” of dynamic information design.
@@
Fix an obedient contract $\sigma$. Construct $\sigma^{*}$ as follows: at any on-path history, replace (relabel) $s$ with the distribution of random signals $\{(\theta|s),a\}$ each with probability $\alpha(s)(a)$, i.e. recommend directly the induced posterior belief and also the recommended action. Clearly this does not affect Receiver’s information or their continuation incentives and hence the realized belief is $\theta|s$ and the recommended action is obeyed. Second, note decreasing transfers as much as possible at actions not induced on path can only weaken incentive compatibility constraints and help support a conjectured equilibrium profile. ∎
@@
The first part is standard, noting that the maximizing set of the static problem is nonempty and that payoffs satisfy a transversality condition (by (BD)). The second follows by setting $u^{\prime}(\mu,a)=\mathcal{U}(\{h^{\xi},\mu,a\})$, so that the choice of future utility promise and payoffs today must be part of an optimum to the sequential formulation, as otherwise a profitable one-shot deviation exists in the sequential problem. ∎
@@
The first statement. For any $(u,\mu_{0})$ and $\{\rho,t,u^{\prime},\alpha\}\in\mathcal{F}(u,\mu_{0})$, and $\varepsilon>0$, let $\tilde{V}(u+\varepsilon,\mu_{0})$ be the value for Sender by choosing at inputs $(u+\varepsilon,\mu_{0})$ the (feasible) tuple $\{\rho,t,u^{\prime}+\frac{\varepsilon}{\delta},\alpha\}$. This tuple strictly increases Receiver utility to meet the higher utility promise without otherwise affecting any IC constraints, so in particular it must do weakly worse than the optimum: $V(u+\varepsilon,\mu_{0})\geq\tilde{V}(u+\varepsilon,\mu_{0})$.
Suppose now the lemma is false, so $V_{+}^{\prime}(u^{\prime}(\mu,a),M\mu)>-k$ for some $u^{\prime}(\mu,a)$ where $\mu\in\text{supp}(\rho)$, $a\in\text{supp}(\alpha)$. By construction, we have
@@
$\tilde{V}(u+\varepsilon,\mu_{0})-V(u,\mu_{0})=\delta\mathbb{E}_{\rho,\alpha}\left[V\left(u^{\prime}(\mu,a)+\frac{\varepsilon}{\delta},M\mu\right)-V(u^{\prime}(\mu,a),M\mu)\right].$
@@
$=\lim\limits_{\varepsilon\to 0_{+}}\frac{\delta}{\varepsilon}\mathbb{E}_{\rho,\alpha}\left[V\left(u^{\prime}(\mu,a)+\frac{\varepsilon}{\delta},M\mu\right)-V(u^{\prime}(\mu,a),M\mu)\right]$
$=\mathbb{E}_{\rho,\alpha}\left[V_{+}^{\prime}(u^{\prime}(\mu,a),M\mu)\right]>-k$
where the second equality interchanges the (finite) expectation and evaluates the right derivative explicitly. The final inequality uses the fact that $V_{+}^{\prime}(\cdot,M\mu)\geq-k$ everywhere with strict inequality (by assumption) for at least one $(\mu,a)\in\text{supp}(\rho,\alpha)$. This contradicts our assumption about $V_{+}^{\prime}(u,\mu_{0})$.
The second statement. Suppose not. Then we can find tuple $\{\rho,t,u^{\prime},\alpha\}\in\mathcal{F}(u,\mu_{0})$ at some $(u,\mu_{0})$ where $t(\mu,a)>0$, $\mu\in\text{supp}(\rho)$, $a\in\text{supp}(\alpha)$, but $V_{+}^{\prime}(u^{\prime}(\mu,a),\mu)>-k$. Recall Receiver’s incentive compatibility constraint when recommended $a$ at $\mu$ takes the form
@@
Consider now the alternative compensation scheme $\{\rho,\tilde{t},\tilde{u},a\}$ which leaves the information and action recommendations unchanged by changing the compensation scheme only at $\mu$: for some $\varepsilon>0$ (restricted to be small enough that $\tilde{t}(\mu)$ remains positive), set $\tilde{t}(\mu,a)=t(\mu,a)-\varepsilon$, and set $\tilde{u}^{\prime}(\mu,a)=u^{\prime}(\mu,a)+\frac{1-\delta}{\delta}\varepsilon$ at all other beliefs $\tilde{t}=t$ and $\tilde{u}^{\prime}=u^{\prime}$. Then by construction
$(1-\delta)\tilde{t}(\mu)+\delta\tilde{u}^{\prime}=(1-\delta)t(\mu)+\delta u^{\prime}$
which implies Receiver’s incentive compatibility constraints are unperturbed under $\{\rho,\tilde{t},\tilde{u}^{\prime},a\}$ (as is their promise-keeping constraint). Hence, $\{\rho,\tilde{t},\tilde{u}^{\prime},a\}$ is feasible at $(u,\mu_{0})$. Whenever belief $\mu$ is realized and $a$ is recommended, however, Sender’s payoff differs by a value of
$(1-\delta)k(t(\mu,a)-\tilde{t}(\mu,a))+\delta(V(\tilde{u}^{\prime}(\mu,a),M\mu)-V(u^{\prime}(\mu,a),M\mu))$
$=(1-\delta)k\varepsilon+\left[\delta V\left(u^{\prime}(\mu,a)+\frac{1-\delta}{\delta}\varepsilon,M\mu\right)-\delta V(u^{\prime}(\mu,a),M\mu)\right].$
@@
$\lim\limits_{\varepsilon\to 0_{+}}\left((1-\delta)k+\delta\frac{V\left(u^{\prime}(\mu,a)+\frac{1-\delta}{\delta}\varepsilon,M\mu\right)-V(u^{\prime}(\mu,a),M\mu)}{\varepsilon}\right)$
$=(1-\delta)k+(1-\delta)V_{+}^{\prime}(u^{\prime}(\mu,a),M\mu).$
@@
Start from $u$-constrained optimal $\sigma$ and define $\bar{\sigma}$ as follows. Let $\rho^{FI}$ be the experiment that reveals the state, and suppose $\bar{\rho}(h^{\xi};\bar{\sigma})=\rho^{FI}$ always. Suppose Sender recommends some action for sure $\bar{\alpha}(\delta_{\theta};\bar{\sigma})=\delta_{a(\theta)}$, where $a(\theta)\in\operatorname*{\arg\!\max}_{a\in A}u(a,\theta)$ is some action that maximizes Receiver’s utility at that state, and which promises Receiver $u(\delta_{\theta},a)=u^{RFI}(\delta_{\theta})$ in the future at all recommended actions (and $\underline{U}(\delta_{\theta})$ otherwise). Suppose moreover that payments never occur on the path of play, $\bar{t}\equiv 0$. Clearly $\{\bar{\rho},\bar{t},\bar{\alpha}\}$ is a feasible contract at any $u\leq u^{RFI}(\mu_{0})$ and gives Receiver a payoff of exactly $u^{RFI}(\mu_{0})$.
Now define the following contract $\sigma^{\prime}$. Still always reveal the state at each history, and let $\alpha^{\prime}$ be defined as follows: for each state $\theta$ and on-path $h^{\xi}$, set $\alpha^{\prime}(h^{\xi};\sigma^{\prime})(\delta_{\theta})=\mathbb{Q}_{\xi}^{\sigma}(\cdot|\theta)\in\Delta(A)$ and let transfers $t^{\prime}$ be such that for each $a^{\prime}\in\text{supp}(\alpha^{\prime}(h^{\xi};\sigma^{\prime})(\delta_{\theta}))$,
@@
$(1-\delta)\max_{a\in A}u(a,\theta)+\delta u^{RFI}(M\delta_{\theta})\geq(1-\delta)u(\tilde{a},\theta)+\delta\underline{U}(M\delta_{\theta})$
@@
We thus now have a profile $\{\bar{\rho},t^{\prime},\alpha^{\prime}\}$ that, by construction, satisfies $\mathbb{Q}_{\xi}^{\sigma}=\mathbb{Q}_{\xi}^{\sigma^{\prime}}$ at every possible history and where the state is fully revealed, i.e. properties (1) and (2). Moreover, it gives (by construction) Sender the same expected payoff absent transfers. How much more must Sender pay Receiver in expectation under $\sigma^{\prime}$? Note by construction additional payments are exactly equal to
$(1-\delta)\sum_{\xi=0}^{\infty}\delta^{\xi}\mathbb{E}_{Q_{\xi}^{\sigma}}\left[\max_{a\in A}u(a,\theta)-u(a^{\prime},\theta)\right]=u^{RFI}(\mu_{0})-\mathcal{U}(\sigma)+\mathcal{T}(\sigma),$
where $\mathcal{T}(\sigma)$ are normalized payments under $\sigma$. Here, we use the fact
@@
$\mathcal{V}(\sigma^{\prime})=\mathcal{V}(\sigma)-k\left(u^{RFI}(\mu_{0})-\mathcal{U}(\sigma)+\mathcal{T}(\sigma)\right)\implies V(\sigma)-V(\sigma^{\prime})\leq k(u^{RFI}(\mu_{0})-\mathcal{U}(\sigma)).$